In an acidic solution with a pH of 3.5, 78% of the
Answer #1

pH = -log [H+]
3.5 = -log [H+]
[H+] = 3.16*10^-4 M

let the acid be written as HA
HA <--------> H+ + A-
C 0 0 (initial)
C-x x x (at equilibrium)

x = [H+] = 3.16*10^-4 M

x = 78% of C
3.16*10^-4 = 0.78*C
C = 4.05*10^-4 M

Ka = x*x / (C-x)
= (3.16*10^-4 )*(3.16*10^-4 ) / (4.05*10^-4 - 3.16*10^-4 )
= 1.12*10^-3

pKa = -log Ka
= -log (1.12*10^-3)
= 2.95